Bucks Ltd is a small and growing company located in Parramatta CBD. The details of the two machines acquired by the company are provided below. The accountant decides to apply different depreciation methods for the machines. 

MachinePurchase dateCostResidual valueUseful life (in years)Depreciation method
Machine 11 January 2016$110,00030,000-Units-of production
Machine 21 January 2016$165,000$15,00015 yearsStraight -line

Machine Purchase date Cost Residual value 30,000 Useful life in years) 1 January 2016 $110,000 Depreciation method Units-of production Straight-line Machine 2 1 January 2016 $165,000 $15,000 15 years 


It is expected that Machine I will produce total of 20,000 units. Actual units for the first three years were: 2,000 units in 2016; 3,500 units in 2017; and 4,500 in 2018


 Required:

 a) Calculate the amount of depreciation expense for the year ended 31 December 2018 for Machine 1.

 b) Calculate the amount of depreciation expense for the year ended 31 December 2018 for Machine 2.

 c) Prepare the journal entry to record depreciation expense on Machine 2 for the year ended 31 December 2018.


The reporting period ends on 31 December. Ignore GST
Answer #1

a) Depreciation expense for machine 1 for the year ended December 31,2018 = (110000-30000/20000)*4500 = 18000

b) Depreciation expense for Machine 2 for the year ended December 31,2018 = (165000-15000)/15 = 10000

c) Journal entry

Date account and explanation Debit Credit
Dec 31,2018 Depreciation expense 10000
Accumulated depreciation-Machine 2 10000
(To record Dep)